Slow​ 'n Steady,​ Inc., has a stock price of $ 28​, will pay a dividend next year of $ 3.05​, and has expected dividend growth of 1.2 % per year. What is your estimate of Slow​ 'n Steady's cost of equity​ capital?
Answer #1

question 1:

As per Gordon Growth Model of Stock Valuation:

P=D1/(Ke-g)

P= price of share. $28
D1= expected next year dividend dividend $3.05
g= Growth rate 1.2% or 0.012
Ke= cost of equity capital

28= 3.05/(Ke-0.012)
Ke= (3.05/28)+0.012
Ke=0.1209
Ke=12.09%

Cost of equity capital of slow n steady Inc is 12.09%